OpenStudy (anonymous):

Hey! How do you find the equation of the conic section for the conditions:hyperbola eccentricity is 2 and foci are at (1.0) and (-1,0)?:)

OpenStudy (anonymous):

ae=1 a=1/2 as e=2 \[e^2=1+(\frac ba)^2\] b/a=sqrt(3) b=sqrt(3)/2 b^2=3/4 a^2=1/4 \[\frac{x^2}{(\frac14)^2}-\frac{y^2}{(\frac34)^2}=1\]

OpenStudy (anonymous):

ohhok..gotcha! thank u so much:)

OpenStudy (anonymous):

ANY1 MORE PROBLEM IN LINEAR ALGEBRA TELL ME

OpenStudy (anonymous):

r u here

OpenStudy (anonymous):

cm on chat box didi

OpenStudy (anonymous):

yes...i got my answer..if u want u can help me on the previous question:) i posted

OpenStudy (anonymous):

ok